# The Ehrhart polynomial of a matroid specializes to the beta invariant

### v1

@@ Abstract.
We show that the linear coefficient of the Ehrhart polynomial of a matroid base polytope evaluated at $t-1$ is equal to, up to normalization, the $\beta$-invariant of the matroid. This yields a lattice-point counting formula for the $\beta$-invariant and establishes a new and unexpected positivity property of Ehrhart polynomials of matroid polytopes.
@@ 1. Introduction
Theorem 1.1 allows us to derive several previously unknown facts about Ehrhart polynomials of matroids and $\beta$-invariants. First, since the $\beta$-invariant is nonnegative for all matroids, the above formula asserts a new positivity property of Ehrhart polynomials of matroids, namely, that the linear coefficient of $\operatorname{ehr}(\mathsf{M},t-1)$ is nonnegative. We point out that Jochemko–Ravichandran [JR22] and, independently, Castillo–Liu [CL21] proved that the linear term of $\operatorname{ehr}(\mathsf{M},t)$, i.e., $\frac{\mathrm{d}}{\mathrm{d}t}\thinspace\operatorname{ehr}(\mathsf{M},t)\big{|}_{t=0}$, is nonnegative. Despite the resemblance, the aforementioned result does not imply nor is implied by ours.
Second, since the valuativity of the Ehrhart polynomial is obvious from the definition, our result gives a simple conceptualization of why the $\beta$-invariant is also a valuative invariant. The valuativity of the $\beta$-invariant was originally established in the work of Speyer [Spe08] and Ardila–Fink–Rincón [AFR10] via a topological reasoning. It is worth noting that, although our result explains the valuativity of $\beta$, this property of $\beta$ plays a crucial role in our proof.
@@ 2. The proof
Recall that a snake matroid is a connected lattice-path matroid whose skew-shape representation is a border strip, i.e., it does not contain a $2\times 2$ square (for more on lattice-path matroids, see [BdMN03] and for additional details on snake matroids, see [KMSRA18, FS24, BKV23, FMP25]). Every skew shape $\lambda/\mu$ comes equipped with a cell poset $P_{\lambda/\mu}$, whose Hasse diagram is obtained from $\lambda/\mu$ by replacing each box with a vertex, connecting adjacent cells, and rotating clockwise by $45$ degrees. The cell posets of skew-shapes corresponding to snake matroids are thus fence posets in the sense of [Sta12, Exercise 3.66]. In Figure 1, we give an example of a skew shape whose lattice-path matroid is a snake, as well as the corresponding fence poset.
@@
Recently, Ferroni and Schröter [FS24] proved that indicator functions of matroid polytopes can be written as linear combinations of indicator functions of snake-like matroids. This result will be a key step in our proofs, so we make it precise below.
@@
[…]
@@
Since each $\mathscr{P}(\mathsf{M}_{i})$ is a $0/1$-polytope, it does not contain lattice points in its relative interior. Using Ehrhart–Macdonald reciprocity (see [BR15, Chapter 4]), this means that $\operatorname{ehr}(\mathsf{M}_{i},-1)=0$. The only exception would be when $\mathscr{P}(\mathsf{M}_{i})$ is a point, but that case is ruled out because it corresponds to $\mathsf{M}_{i}$ being a direct sum of loops and coloops. Since $s\geq 2$ and each factor $\operatorname{ehr}(\mathsf{M}_{i},t)$ vanishes at $t=-1$, we have that $-1$ is a root of $\operatorname{ehr}(\mathsf{M},t)$ of multiplicity at least $2$, which yields the desired equality. ∎
@@
$[t^{1}]\operatorname{ehr}(\mathsf{M},t-1)=\sum_{i=1}^{m}a_{i}\thinspace[t^{1}]\operatorname{ehr}(\mathsf{S}_{i},t-1).$
By Lemma 2.4, if $\mathsf{S}_{i}$ is disconnected then $[t^{1}]\operatorname{ehr}(\mathsf{S}_{i},t-1)=0$, so we can restrict the sum to (connected) snake matroids.
@@
[…]
@@
Theorem 1.1 is false if we drop the condition that the matroid is both loopless and coloopless. For example, let $\mathsf{M}=\mathsf{U}_{2,4}\oplus\mathsf{U}_{1,1}$ be the direct sum of two uniform matroids. The $\beta$-invariant of $\mathsf{M}$ is 00 since $\mathsf{M}$ is disconnected, but the base polytope of $\mathsf{M}$ is integrally equivalent to the base polytope of $\mathsf{U}_{2,4}$ and thus the linear term of $\operatorname{ehr}(\mathsf{M},t-1)$ coincides with the linear term of $\operatorname{ehr}(\mathsf{U}_{2,4},t-1)$, which is equal to $\frac{1}{3}$.
@@
It follows from [FSS24, Proposition 12.1] that whenever the matroid $\mathsf{M}$ has rank $k$ and size $n=2k$, we have $\omega(\mathsf{M})=(-1)^{c(\mathsf{M})}\cdot\operatorname{ehr}(\mathsf{M},-2)$, where $c(\mathsf{M})$ denotes the number of connected components of $\mathsf{M}$. Therefore, in this special case where $n=2k$, $\omega(\mathsf{M})$ is also a specialization of the Ehrhart polynomial of $\mathsf{M}$. However, if $n\neq 2k$, the last assertion is no longer true. We have verified by an exhaustive computation that the $\omega$-invariant of matroids on $7$ elements of rank $3$ is not a linear specialization of the Ehrhart polynomial: we found a collection of (connected) matroids of this rank and size whose Ehrhart polynomials satisfy a linear relation that is not satisfied by the corresponding $\omega$-invariants.
@@ Acknowledgments
This article resulted from a SQuaRE at the American Institute for Mathematics (AIM). The authors thank AIM for providing a very supportive and mathematically rich environment. Anastasia Chavez is partially supported by NSF grant DMS 2332342. Luis Ferroni is a member at the Institute for Advanced Study, funded by the Minerva Research Foundation. Fu Liu is partially supported by an NSF grant #2153897-0.

### v2

@@ Abstract.
We show that the linear coefficient of the Ehrhart polynomial of a matroid base polytope evaluated at t1t-1 is equal to, up to normalization, the $\beta$-invariant of the matroid. This yields a lattice-point counting formula for the $\beta$-invariant and establishes a new and unexpected positivity property of Ehrhart polynomials of matroid polytopes.
@@ 1. Introduction
Theorem 1.1 allows us to derive several previously unknown facts about Ehrhart polynomials and $\beta$-invariants of matroids. First, since the $\beta$-invariant is nonnegative for all matroids, the above formula asserts a new positivity property of Ehrhart polynomials of matroids, namely, that the linear coefficient of $\operatorname{ehr}(\mathsf{M},t-1)$ is nonnegative. We point out that Jochemko–Ravichandran [JR22] and, independently, Castillo–Liu [CL21] proved that the linear term of $\operatorname{ehr}(\mathsf{M},t)$, i.e., $\frac{\mathrm{d}}{\mathrm{d}t}\thinspace\operatorname{ehr}(\mathsf{M},t)\big|_{t=0}$, is nonnegative. Despite the resemblance, the aforementioned result does not imply nor is implied by ours.
Second, since the valuativity of the Ehrhart polynomial is obvious from the definition, our result gives a simple conceptualization of why the $\beta$-invariant is also a valuative invariant. The valuativity of the $\beta$-invariant was originally established in the work of Speyer [Spe08] and Ardila–Fink–Rincón [AFR10] via topological reasoning. It is worth noting that, although our result explains the valuativity of $\beta$, this property of $\beta$ plays a crucial role in our proof.
@@ 2. The proof
We make a brief recapitulation about lattice-path matroids. Fix two nonnegative integers $k\leq n$. We look at lattice paths in the plane $\mathbb{R}^{2}$ that start at $(0,0)$, end at (nk,k)(n-k,k), and are composed of exactly $n$ unit steps, each being either +(1,0)+(1,0) or +(0,1)+(0,1).
Each lattice path can be represented as a word of length $n$ in the alphabet $\{N,E\}$, where the $i$-th letter is $N$ if the $i$-th step is +(0,1)+(0,1) (“north”) and $E$ if it is +(1,0)+(1,0) (“east”). See Figure 1 for an example of two lattice paths from $(0,0)$ to $(5,4)$.
Figure 1. The paths $L=\text{EEENENNEN}$ and $U=\text{NEENENNEE}$ for $k=4$ and $n=9$. The region between them is shaded.
Alternatively, one may encode a path $P$ by the subset $s(P)\subseteq[n]$ consisting of the indices of its north steps. For the two paths in Figure 1, we have $s(L)=\{4,6,7,9\}$ and $s(U)=\{1,4,6,7\}$.
For fixed integers $k$ and $n$, let $L$ and $U$ be two lattice paths of this type. We say that $L$ lies below $U$ if, for every $m=1,\ldots,n$,
$\bigl|s(L)\cap[m]\bigr|\leq\bigl|s(U)\cap[m]\bigr|\,.$ (1)
In such case, we write $L\leq U$. In Figure 1, for example, $L\leq U$. It follows from [BdMN03, Theorem 3.3] that for every pair of lattice paths $L\leq U$, the collection of sets $\{s(P):L\leq P\leq U\}$ defines the family of bases of a matroid on $[n]$ of rank $k$. Matroids arising in this way are called lattice-path matroids. The skew shape that lies between $L$ and $U$ is often called the skew-shape representation of a lattice-path matroid. For example, in Figure 2 (on the left) we see the skew-shape representation of the lattice-path matroid of rank $4$ on $n=9$ elements, having $s(L)=\{4,6,7,9\}$ and $s(U)=\{1,4,6,7\}$. It is a well-known fact that a lattice-path matroid is connected if and only if the paths $L$ and $U$ intersect only at $(0,0)$ and (nk,k)(n-k,k) (see [BdMN03, Theorem 3.6]).
A snake matroid is a connected lattice-path matroid whose skew-shape representation is a border strip, i.e., it does not contain a $2\times 2$ square (for additional details on snake matroids, see [KMSRA18, FS24, BKV23, FMP25]). Every skew shape $\lambda/\mu$ comes equipped with a cell poset $P_{\lambda/\mu}$, whose Hasse diagram is obtained from $\lambda/\mu$ by replacing each box with a vertex, connecting adjacent cells, and rotating clockwise by $45$ degrees. The cell posets of skew shapes corresponding to snake matroids are thus fence posets in the sense of [Sta12, Exercise 3.66]. In Figure 2, we give an example of a skew shape whose lattice-path matroid is a snake, as well as the corresponding fence poset.
@@
Recently, Ferroni and Schröter [FS24] proved that indicator functions of matroid polytopes can be written as linear combinations of indicator functions of direct sums of snake matroids, loops, and coloops. This result will be a key step in our proofs, so we make it precise below.
@@
    We have $\amathbb{1}_{\mathscr{P}(\mathsf{M})}=\sum_{i=1}^{m}a_{i}\,\amathbb{1}_{\mathscr{P}(\mathsf{S}_{i})}$.
@@
Since each $\mathscr{P}(\mathsf{M}_{i})$ is a $0/1$-polytope, it does not contain lattice points in its relative interior. Using Ehrhart–Macdonald reciprocity (see [BR15, Chapter 4]), this means that $\operatorname{ehr}(\mathsf{M}_{i},-1)=0$. The only exception would be when $\mathscr{P}(\mathsf{M}_{i})$ is a point, but that case is ruled out because it corresponds to $\mathsf{M}_{i}$ being a direct sum of loops and coloops. Since $s\geq 2$ and each factor $\operatorname{ehr}(\mathsf{M}_{i},t)$ vanishes at t=1t=-1, we have that 1-1 is a root of $\operatorname{ehr}(\mathsf{M},t)$ of multiplicity at least $2$, which yields the desired equality. ∎
@@
$t^{1}]\operatorname{ehr}(\mathsf{M},t-1)=\sum_{i=1}^{m}a_{i}\thinspace[t^{1}]\operatorname{ehr}(\mathsf{S}_{i},t-1).$
By Lemma 2.4, if $\mathsf{S}_{i}$ is disconnected then $t^{1}]\operatorname{ehr}(\mathsf{S}_{i},t-1)=0$, so we can restrict the sum to (connected) snake matroids.
@@
$[t^{1}]\operatorname{ehr}(\mathsf{M},t-1)$ $=\sum_{\text{$\mathsf{S}_{i}$ connected}}a_{i}\thinspace[t^{1}]\operatorname{ehr}(\mathsf{S}_{i},t-1)$
$=\sum_{\text{$\mathsf{S}_{i}$ connected}}a_{i}\thinspace\frac{1}{(n-1)\binom{n-2}{k-1}}$
@@
Theorem 1.1 is false if we drop the condition that the matroid is both loopless and coloopless. For example, let $\mathsf{M}=\mathsf{U}_{2,4}\oplus\mathsf{U}_{1,1}$ be the direct sum of two uniform matroids. The $\beta$-invariant of $\mathsf{M}$ is 0 since $\mathsf{M}$ is disconnected, but the base polytope of $\mathsf{M}$ is integrally equivalent to the base polytope of $\mathsf{U}_{2,4}$ and thus the linear term of $\operatorname{ehr}(\mathsf{M},t-1)$ coincides with the linear term of $\operatorname{ehr}(\mathsf{U}_{2,4},t-1)$, which is equal to $\frac{1}{3}$.
Example 2.6
Let us consider the uniform matroid $\mathsf{U}_{3,8}$. The $\beta$-invariant of the uniform matroid $\mathsf{U}_{k,n}$ equals $\binom{n-2}{k-1}$ (see, for example, [Cra67, Proposition 7]), so we have $\beta(\mathsf{U}_{3,8})=\binom{6}{2}=15$. Theorem 1.1 says that
$t^{1}]\operatorname{ehr}(\mathsf{U}_{3,8},t-1)=\frac{15}{7\cdot\binom{6}{2}}=\frac{1}{7}.$
This can be verified directly, for instance using closed formulas for the Ehrhart polynomial of a uniform matroid (see for example [Kat05, Fer21]). In particular, [Fer21, Proposition 2.1] gives
$\operatorname{ehr}(\mathsf{U}_{3,8},t-1)=\frac{397}{1680}t^{7}+\frac{49}{144}t^{6}+\frac{1}{48}t^{5}+\frac{19}{144}t^{4}+\frac{1}{10}t^{3}+\frac{1}{36}t^{2}+\frac{1}{7}t,$
where the linear terms matches the expected value.
@@
It follows from [FSS24, Proposition 12.1] that whenever the matroid $\mathsf{M}$ has rank $k$ and size $n=2k$, we have $\omega(\mathsf{M})=(-1)^{c(\mathsf{M})}\cdot\operatorname{ehr}(\mathsf{M},-2)$, where $c(\mathsf{M})$ denotes the number of connected components of $\mathsf{M}$. Therefore, in this special case where $n=2k$, $\omega(\mathsf{M})$ is also a specialization of the Ehrhart polynomial of $\mathsf{M}$. However, if $n\neq 2k$, the last assertion is no longer true. We have verified by an exhaustive computation that the $\omega$-invariant of matroids on $7$ elements of rank $3$ is not a linear specialization of the Ehrhart polynomial: we found a collection of connected matroids of this rank and size whose Ehrhart polynomials satisfy a linear relation that is not satisfied by the corresponding $\omega$-invariants.
@@ Acknowledgments
This article resulted from a SQuaRE at the American Institute for Mathematics (AIM). The authors thank AIM for providing a very supportive and mathematically rich environment, and are grateful to two anonymous reviewers for their helpful comments and suggestions that improved this article. Anastasia Chavez is partially supported by NSF grant DMS 2332342. Luis Ferroni was a member at the Institute for Advanced Study, funded by the Minerva Research Foundation. Fu Liu is partially supported by an NSF grant #2153897-0.